OpenStudy (anonymous):

given the points (2,3) and (6,8) write the point slope form of the equation A.y-3=x-2 B.y-3=5/4 (x-2) C.y-8=4/5(x-6) ??

OpenStudy (anonymous):

First find the slope using (y2-y1)/(x2-x1)

OpenStudy (anonymous):

to find the slope we use the equation y2-y1/x2-x1 8-3/6-2=5/4, so 5/4 is our slope. Next we use the point slope form equation y-y1=m(x-x1), and plug in our points and slope to get y-3=5/4(x-2)

OpenStudy (anonymous):

thank you!!